How many grams of solid zinc is needed to react with 25.0 mL of a 1.5 M solution of hydrochloric acid (HCl)?
1
Expert's answer
2018-03-19T08:44:27-0400
V(HCl)=25 mL
(HCl)=1.5M
.m(Zn)=?
We write the equation between zinc and hydrochloric acid
Zn + 2HCl=ZnC + 2
Find the amount of substance of HCl
.n(HCl)= ×V=1.5 × 0.025=0.0375 mol
.n(Zn) = n(HCl)= ×0.0375 mol=0.01875 mol
Find the mass of Zn
.m(Zn)= n(Zn)×Ar(Zn)=0.0187565.4=1.23 g